3x+5y=-11
x-2y=11

Answer by themathtutor2009(81)

3x+5y=-11
x-2y=11



3x+5y=-11
-3x+6y=-33

11y=-44
y=-4

x-2y=11
x-2(-4)=11
x+8=11
x=3

Answers are x=3, y=-4